) Determine a unit vector perpendicular to the plane of A = ˆi − 2ˆj + kˆ


r


and B = 2ˆi + 3ˆj − kˆ


r


Grade:

1 Answers

APURV GOEL
12 years ago

Assume the vector to be ai + bj + ck

as it is perpendicular to other two vectors dot product must be 0

so we get a - 2b + c = 0 by taking dot prod. with first one

and 2a + 3b - c = 0 by taking dot prod. with second one

and it is a unit vector this means a^2 + b^2 + c^2 = 1

you have three variables and three equations solve to get the answer.

Please approve if u got what u were looking for.